OpenStudy (anonymous):

3/4 - 3/10 = ? need in simplest form

OpenStudy (lieutenantgeneral):

9/20

OpenStudy (anonymous):

find a common denominator. The least common multiple is 20. So multiply 3/4 with 5 and 3/10 with 2. So it would become 15/20 - 6/20. 15- 6= 9. so the answer is 9/20

OpenStudy (anonymous):

ok 3/4 - 3/10=0.45 and 0.45 in simplest form is 45/100 which changes to 9/20

OpenStudy (anonymous):

4/7 - 2/5 = ? simplest form

OpenStudy (anonymous):

find a common denominator once again. The least common denominator is 35. So multiply 4/7 with 5 and 2/5 with 7. This becomes 20/35 - 14/ 35. So the answer is 6/ 35.

OpenStudy (anonymous):

4/7 - 2/5 =0.17142857142

OpenStudy (anonymous):

nice use of a calculator lol

OpenStudy (anonymous):

you should change the denominator so their the same or you could type it in www.google.com or www.bing.com

OpenStudy (anonymous):

not a calculator im fond of checking my answers before helping someone thank u very much @Princesslord1000

OpenStudy (anonymous):

You try finding out how to do it o another website? that's what I usually do!!!!

OpenStudy (anonymous):

ya i have it usually helps

OpenStudy (anonymous):

@Starr_DynastyT Then do it again and see what you find out

OpenStudy (anonymous):

ya i did already i got the answer trying to help out @NatPanda by explaining it